Assume the following cost data relate to the decision to produce

Total units expected to be produced or bought from the market = 10,000

Total Costs

Unit Cost

Direct materials

20,000

2.00

Direct labor

25,000

2.50

Variable overhead

15,000

1.50

Fixed overhead (non-avoidable)

24000

2.40

Fixed overhead (avoidable)

26,000

2.60

Purchase cost

86,000

Should the company acquire the product form the market?

a.

Yes

b.

No

c.

Indifferent to to make or to buy

d.

Yes if the market price per unit less than $4.50
Answer #1

Answer

Total cost produce 10,000 units

= 20,000 + 25,000 + 15,000 + 24,000 + 26,000 = $110,00

Out of $110,000 production cost non avoidable cost is 24,000

Avoidable cost = 110,000 - 24,000 = $86,000

Purchase cost is $86,000

Non avoidable cost are irrelevant in make or buy decisions. Managerial decisions are not affected to non avoidable cost. The non avoidable cost of $24,000 incurred if company make or buy the product.

So the production cost and purchase cost are same ($86,000). If the company make or buy the product , total cost is same $110,000 (86,000 + 24,000).

The company is indifferent to make or buy decision.

Option C is CORRECT indifferent to to make or buy.

Option A and B is INCORRECT because the making cost and purchase cost are same $86,000. Its indifferent to make or buy decision.

Option d is INCORRECT because if the market price per unit less than $4.5 , but the purchase cost is $86,000. Purchase may include transportation charges , import duty , taxes etc. Purchase cost not include only market price.

Option C is CORRECT.

If YOU are like the answer please upvote.